OpenStudy (anonymous):

Help with systems of solving equation with three variables (I will post the question)

OpenStudy (anonymous):

2x + y - 2z = 3 x- y + z = 3 x+ 3y - 3z = 3

OpenStudy (anonymous):

I tried using elimination method by first taking equation 1 and 2, then equation 2 and 3 I got these equations as a result: 3x - z = 6 -2x - 6z = -6

OpenStudy (anonymous):

But when i tried to use elimination for those two equations I ended up with a fraction x= 21/10

OpenStudy (thomas5267):

Nope. The answers are all in nice integers.

OpenStudy (anonymous):

Could someone check my work and tell me where I went wrong?

OpenStudy (thomas5267):

Your second equation is wrong I think. 3 times the second equation is then added to the third one not subtracted.

OpenStudy (thomas5267):

x - y (yes that is a minus y) + z = 3

OpenStudy (anonymous):

This is what I did for the second and thrid equation -3x - 3y -3z = -9 (i multplied this one by -3) + x + 3y - 3z =3 and I got -2x -6z = -6 when I added them

OpenStudy (thomas5267):

\[ x{\color{red}-}y+z=3\\ \]